OpenStudy (anonymous):

***WILL GIVE MEDAL TO WHOEVER HELPS*** Question: Sherri has 23 pieces of jewelry to sell. She sells the bracelets for $2 and the necklaces for $3, and earns a total of $55. If the bracelets are represented by x and the necklaces are represented by y, which of the following systems of equations can be used to calculate the number of bracelets and necklaces sold? Answer choices: x + y = 23, 3x + 2y = 55 x + y = 55, 3x + 2y = 23 x + y = 23, 2x + 3y = 55 x + y = 55, 2x + 3y = 23

OpenStudy (texaschic101):

x = bracelets y = necklaces there are 23 of them in total (necklaces + bracelets) x + y = 23 bracelets are $2 each and necklaces are $3 each...totaling $55 2x + 3y = 55

OpenStudy (anonymous):

texas thanks

OpenStudy (texaschic101):

sure thing :)

OpenStudy (anonymous):

can u help with another

OpenStudy (texaschic101):

possibly...what ya got ??

OpenStudy (anonymous):

i got the hang of it but not quite sure

OpenStudy (anonymous):

Solve the following system of equations by substitution and select the correct answer below: 6x − 4y = 36 2x − 8y = 32 x = 4, y = 3 x = 4, y = −3 x = −4, y = 3 x = −4, y = −3

OpenStudy (anonymous):

3rd choice was my answer

OpenStudy (texaschic101):

2x - 8y = 32 --- reduce by dividing by 2 x - 4y = 16 x = 4y + 16 now sub 4y + 16 in for x back into the other equation 6x - 4y = 36...this can be reduced as well 3x - 2y = 18 now sub 3(4y + 16) - 2y = 18 12y + 48 - 2y = 18 subtract 48 12y - 2y = 18 - 48 10y = - 30 y = -3 x = 4y + 16 x = 4(-3) + 16 x = -12 + 16 x = 4 check... x = 4y + 16 4 = 4(-3) + 16 4 = -12 + 16 4 = 4 (correct) answer is (4,-3)

OpenStudy (anonymous):

thank u now i understand it :)

OpenStudy (anonymous):

just need this last one and im done youve been such a good help

OpenStudy (anonymous):

Is this right? What are the solutions to the following system of equations? Select the correct answer below. y = x2 − 4x + 8 4x + y = 12 (2, 20) and (−2, 4) (2, 4) and (−2, 20) (−2, 2) and (4, 20) (20, −2) and (4, 2) <------- my answer

OpenStudy (texaschic101):

4x + y = 12 y = -4x + 12 x^2 - 4x + 8= -4x + 12 x^2 + 8 = 12 x^2 = 4 x = sq rt 4 = 2 and -2 y = -4x + 12 y = -4 (+-) 2 + 12 y = -4 (+2) + 12 y = -8 + 12 y = 4 y = -4 (-2) + 12 y = 8 + 12 y = 20 so y = 4 and 20 so your answers are correct

OpenStudy (anonymous):

thank u texas appreciate all the help

OpenStudy (texaschic101):

sorry it took so long....I hope I helped

OpenStudy (anonymous):

helped a lot